# Functional differential equations driven by càdlàg rough paths

### v2

@@ 2.3. Continuity of the Itô–Lyons map
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[t_{i},t_{i+1}]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[t_{i},t_{i+1}]}$
$\quad\leq\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[t_{i},t_{i+1})}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[t_{i},t_{i+1})}$
$\qquad+|Y^{\prime}_{t_{i+1}-,t_{i+1}}-\tilde{Y}^{\prime}_{t_{i+1}-,t_{i+1}}|+|R^{Y}_{t_{i+1}-,t_{i+1}}-R^{\tilde{Y}}_{t_{i+1}-,t_{i+1}}|$
(2.9) $\quad\lesssim|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\tilde{X},p,[t_{i},t_{i+1}]}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{i},t_{i+1}]},$
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\tilde{\mathbf{X}}\|_{p}$, and not on the index $i$. Here, we used Assumption (2.1) and the estimate (2.8) to derive that
$|(R^{Y}_{t_{i+1}-,t_{i+1}}-R^{\tilde{Y}}_{t_{i+1}-,t_{i+1}})-(R^{y}_{t_{i+1}-,t_{i+1}}-R^{\tilde{y}}_{t_{i+1}-,t_{i+1}})|$
$\quad\leq|F^{\prime}_{t_{i+1}-}(Y,Y^{\prime})\mathbb{X}_{t_{i+1}-,t_{i+1}}-\tilde{F}^{\prime}_{t_{i+1}-}(\tilde{Y},\tilde{Y}^{\prime})\tilde{\mathbb{X}}_{t_{i+1}-,t_{i+1}}|$
$\quad\leq|F^{\prime}_{t_{i+1}-}(Y,Y^{\prime})-\tilde{F}^{\prime}_{t_{i+1}-}(\tilde{Y},\tilde{Y}^{\prime})||\mathbb{X}_{t_{i+1}-,t_{i+1}}|+|\tilde{F}_{t_{i+1}-}^{\prime}(\tilde{Y},\tilde{Y}^{\prime})||\mathbb{X}_{t_{i+1-},t_{i+1}}-\tilde{\mathbb{X}}_{t_{i+1}-,t_{i+1}}|$
$\quad\leq(|F^{\prime}_{t_{i}}(Y,Y^{\prime})-\tilde{F}^{\prime}_{t_{i}}(\tilde{Y},\tilde{Y})|+\|F^{\prime}(Y,Y^{\prime})-\tilde{F}^{\prime}(\tilde{Y},\tilde{Y}^{\prime})\|_{p,[t_{i},t_{i+1})})\|\mathbb{X}\|_{\frac{p}{2},[t_{i},t_{i+1}]}$
$\qquad+(|\tilde{F}^{\prime}_{t_{i}}(\tilde{Y},\tilde{Y}^{\prime})|+\|\tilde{F}^{\prime}(\tilde{Y},\tilde{Y}^{\prime})\|_{p,[t_{i},t_{i+1})})\|\mathbb{X}-\tilde{\mathbb{X}}\|_{\frac{p}{2},[t_{i},t_{i+1}]}$
$\quad\lesssim|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p,[t_{i},t_{i+1})}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{i},t_{i+1}]},$
@@
$|(Y^{\prime}_{t_{i+1}-,t_{i+1}}-\tilde{Y}^{\prime}_{t_{i+1}-,t_{i+1}})-(y^{\prime}_{t_{i+1}-,t_{i+1}}-\tilde{y}^{\prime}_{t_{i+1}-,t_{i+1}})|$
$\quad=|F_{t_{i+1}-,t_{i+1}}(Y)-\tilde{F}_{t_{i+1}-,t_{i+1}}(\tilde{Y})|$
$\quad\leq|(F-\tilde{F})_{t_{i+1}-,t_{i+1}}(Y)|+|\tilde{F}_{t_{i+1}-,t_{i+1}}(Y)-\tilde{F}_{t_{i+1}-,t_{i+1}}(\tilde{Y})|$
$\quad\leq C_{F-\tilde{F}}(1+(|Y^{\prime}_{t_{i}}|+\|Y^{\prime}\|_{p,[t_{i},t_{i+1}]})\|X\|_{p,[t_{i},t_{i+1}]}+\|R^{Y}\|_{\frac{p}{2},[t_{i},t_{i+1}]})$
$\qquad+C_{F,K,X,\tilde{X}}(|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+\|Y-\tilde{Y}\|_{p,[t_{i},t_{i+1})}+|Y_{t_{i+1}-,t_{i+1}}-\tilde{Y}_{t_{i+1}-,t_{i+1}}|)$
$\quad\lesssim|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p,[t_{i},t_{i+1})}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{i},t_{i+1}]},$
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$ and $\|\mathbf{X}\|_{p}\allowbreak\vee\|\tilde{\mathbf{X}}\|_{p}$, and not on the index $i$.
Now, we need to control the term $|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$. For this, we note that
$|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\quad\leq|Y_{t_{i-1}}-\tilde{Y}_{t_{i-1}}|+|F_{t_{i-1}}(Y)-\tilde{F}_{t_{i-1}}(\tilde{Y})|+|y^{\prime}_{t_{i-1}}-\tilde{y}^{\prime}_{t_{i-1}}|$
$\qquad+\|Y-\tilde{Y}\|_{p,[t_{i-1},t_{i}]}+\|F(Y)-\tilde{F}(\tilde{Y})\|_{p,[t_{i-1},t_{i}]}+\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{i-1},t_{i}]}$
$\quad\lesssim|Y_{t_{i-1}}-\tilde{Y}_{t_{i-1}}|+|F_{t_{i-1}}(Y)-\tilde{F}_{t_{i-1}}(\tilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\tilde{X},p,[t_{i-1},t_{i}]}$
$\qquad+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{i-1},t_{i}]}$
$\qquad+\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[t_{i-1},t_{i}]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[t_{i-1},t_{i}]},$
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\tilde{X}}$, $K$, $\|X\|_{p}$ and $\|\tilde{X}\|_{p}$, and not on the index $i$; thus applying the estimate (2.3) for $i-1$ gives that
$|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\quad\lesssim|Y_{t_{i-1}}-\tilde{Y}_{t_{i-1}}|+|F_{t_{i-1}}(Y)-\tilde{F}_{t_{i-1}}(\tilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\tilde{X},p,[t_{i-1},t_{i}]}$
$\qquad+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{i-1},t_{i}]}.$
@@
$|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\tilde{F}_{0}(\tilde{y})|+|y^{\prime}_{0}-\tilde{y}^{\prime}_{0}|$
$\qquad+\sum_{j=0}^{i-1}\Big{(}\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{j},t_{j+1}]}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{j},t_{j+1}]}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{j},t_{j+1}]}\Big{)},$
@@
$(|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|)^{p}$
$\quad\lesssim(|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\tilde{F}_{0}(\tilde{y})|+|y^{\prime}_{0}-\tilde{y}^{\prime}_{0}|+NC_{F-\tilde{F}})^{p}$
$\qquad+\sum_{j=0}^{i-1}\Big{(}\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{j},t_{j+1}]}^{p}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{j},t_{j+1}]}^{p}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p,[t_{j},t_{j+1}]}^{p}\Big{)}.$
@@
$|Y_{t_{i}}-\tilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\tilde{F}_{t_{i}}(\tilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\tilde{F}_{0}(\tilde{y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\tilde{X},p}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p},$
@@
$\|Y^{\prime}-\tilde{Y}^{\prime}\|_{p,[t_{i},t_{i+1}]}+\|R^{Y}-R^{\tilde{Y}}\|_{\frac{p}{2},[t_{i},t_{i+1}]}$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\tilde{F}_{0}(\tilde{y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\tilde{X},p}+C_{F-\tilde{F}}+\|\mathbf{X};\tilde{\mathbf{X}}\|_{p}.$

### v3

@@ 2.3. Continuity of the Itô–Lyons map
(2.7) $\begin{split}&\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[t_{i},t_{i+1}]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[t_{i},t_{i+1}]}\\ &\quad\leq\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[t_{i},t_{i+1})}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[t_{i},t_{i+1})}\\ &\qquad+|Y^{\prime}_{t_{i+1}-,t_{i+1}}-\widetilde{Y}^{\prime}_{t_{i+1}-,t_{i+1}}|+|R^{Y}_{t_{i+1}-,t_{i+1}}-R^{\widetilde{Y}}_{t_{i+1}-,t_{i+1}}|\\ &\quad\lesssim|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\widetilde{X},p,[t_{i},t_{i+1}]}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[t_{i},t_{i+1}]},\end{split}$
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\widetilde{X}}$, $K$, $\|\mathbf{X}\|_{p}$ and $\|\widetilde{\mathbf{X}}\|_{p}$, and not on the index $i$. Here, we used Assumption 2.4 and the estimate (2.6) to derive that
$|(R^{Y}_{t_{i+1}-,t_{i+1}}-R^{\widetilde{Y}}_{t_{i+1}-,t_{i+1}})-(R^{y}_{t_{i+1}-,t_{i+1}}-R^{\tilde{y}}_{t_{i+1}-,t_{i+1}})|$
$\quad\leq|F^{\prime}_{t_{i+1}-}(Y,Y^{\prime})\mathbb{X}_{t_{i+1}-,t_{i+1}}-\widetilde{F}^{\prime}_{t_{i+1}-}(\widetilde{Y},\widetilde{Y}^{\prime})\widetilde{\mathbb{X}}_{t_{i+1}-,t_{i+1}}|$
$\quad\leq|F^{\prime}_{t_{i+1}-}(Y,Y^{\prime})-\widetilde{F}^{\prime}_{t_{i+1}-}(\widetilde{Y},\widetilde{Y}^{\prime})||\mathbb{X}_{t_{i+1}-,t_{i+1}}|+|\widetilde{F}_{t_{i+1}-}^{\prime}(\widetilde{Y},\widetilde{Y}^{\prime})||\mathbb{X}_{t_{i+1-},t_{i+1}}-\widetilde{\mathbb{X}}_{t_{i+1}-,t_{i+1}}|$
$\quad\leq(|F^{\prime}_{t_{i}}(Y,Y^{\prime})-\widetilde{F}^{\prime}_{t_{i}}(\widetilde{Y},\widetilde{Y})|+\|F^{\prime}(Y,Y^{\prime})-\widetilde{F}^{\prime}(\widetilde{Y},\widetilde{Y}^{\prime})\|_{p,[t_{i},t_{i+1})})\|\mathbb{X}\|_{\frac{p}{2},[t_{i},t_{i+1}]}$
$\qquad+(|\widetilde{F}^{\prime}_{t_{i}}(\widetilde{Y},\widetilde{Y}^{\prime})|+\|\widetilde{F}^{\prime}(\widetilde{Y},\widetilde{Y}^{\prime})\|_{p,[t_{i},t_{i+1})})\|\mathbb{X}-\widetilde{\mathbb{X}}\|_{\frac{p}{2},[t_{i},t_{i+1}]}$
$\quad\lesssim|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p,[t_{i},t_{i+1})}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[t_{i},t_{i+1}]},$
@@
$|(Y^{\prime}_{t_{i+1}-,t_{i+1}}-\widetilde{Y}^{\prime}_{t_{i+1}-,t_{i+1}})-(y^{\prime}_{t_{i+1}-,t_{i+1}}-\tilde{y}^{\prime}_{t_{i+1}-,t_{i+1}})|$
$\quad=|F_{t_{i+1}-,t_{i+1}}(Y)-\widetilde{F}_{t_{i+1}-,t_{i+1}}(\widetilde{Y})|$
$\quad\leq|(F-\widetilde{F})_{t_{i+1}-,t_{i+1}}(Y)|+|\widetilde{F}_{t_{i+1}-,t_{i+1}}(Y)-\widetilde{F}_{t_{i+1}-,t_{i+1}}(\widetilde{Y})|$
$\quad\leq C_{F-\widetilde{F}}(1+(|Y^{\prime}_{t_{i}}|+\|Y^{\prime}\|_{p,[t_{i},t_{i+1}]})\|X\|_{p,[t_{i},t_{i+1}]}+\|R^{Y}\|_{\frac{p}{2},[t_{i},t_{i+1}]})$
$\qquad+C_{F,K,X,\widetilde{X}}(|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+\|Y-\widetilde{Y}\|_{p,[t_{i},t_{i+1})}+|Y_{t_{i+1}-,t_{i+1}}-\widetilde{Y}_{t_{i+1}-,t_{i+1}}|)$
$\quad\lesssim|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,p,[t_{i},t_{i+1})}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[t_{i},t_{i+1}]},$
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\widetilde{X}}$, $K$ and $\|\mathbf{X}\|_{p}\allowbreak\vee\|\widetilde{\mathbf{X}}\|_{p}$, and not on the index $i$.
Now, we need to control the term $|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$. For this, we note that
$|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\quad\leq|Y_{t_{i-1}}-\widetilde{Y}_{t_{i-1}}|+|F_{t_{i-1}}(Y)-\widetilde{F}_{t_{i-1}}(\widetilde{Y})|+|y^{\prime}_{t_{i-1}}-\tilde{y}^{\prime}_{t_{i-1}}|$
$\qquad+\|Y-\widetilde{Y}\|_{p,[t_{i-1},t_{i}]}+\|F(Y)-\widetilde{F}(\widetilde{Y})\|_{p,[t_{i-1},t_{i}]}+\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{i-1},t_{i}]}$
$\quad\lesssim|Y_{t_{i-1}}-\widetilde{Y}_{t_{i-1}}|+|F_{t_{i-1}}(Y)-\widetilde{F}_{t_{i-1}}(\widetilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\widetilde{X},p,[t_{i-1},t_{i}]}$
$\qquad+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[t_{i-1},t_{i}]}$
$\qquad+\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[t_{i-1},t_{i}]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[t_{i-1},t_{i}]},$
where the implicit multiplicative constant depends on $p$, $C_{F}$, $C_{F,K,X,\widetilde{X}}$, $K$, $\|X\|_{p}$ and $\|\widetilde{X}\|_{p}$, and not on the index $i$; thus applying the estimate (2.7) for $i-1$ gives that
$|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\quad\lesssim|Y_{t_{i-1}}-\widetilde{Y}_{t_{i-1}}|+|F_{t_{i-1}}(Y)-\widetilde{F}_{t_{i-1}}(\widetilde{Y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\widetilde{X},p,[t_{i-1},t_{i}]}$
$\qquad+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[t_{i-1},t_{i}]}.$
@@
$|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\widetilde{F}_{0}(\tilde{y})|+|y^{\prime}_{0}-\tilde{y}^{\prime}_{0}|$
$\qquad+\sum_{j=0}^{i-1}\Big{(}\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{j},t_{j+1}]}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{j},t_{j+1}]}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[t_{j},t_{j+1}]}\Big{)},$
@@
$(|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|)^{p}$
$\quad\lesssim(|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\widetilde{F}_{0}(\tilde{y})|+|y^{\prime}_{0}-\tilde{y}^{\prime}_{0}|+NC_{F-\widetilde{F}})^{p}$
$\qquad+\sum_{j=0}^{i-1}\Big{(}\|y^{\prime}-\tilde{y}^{\prime}\|_{p,[t_{j},t_{j+1}]}^{p}+\|R^{y}-R^{\tilde{y}}\|_{\frac{p}{2},[t_{j},t_{j+1}]}^{p}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p,[t_{j},t_{j+1}]}^{p}\Big{)}.$
@@
$|Y_{t_{i}}-\widetilde{Y}_{t_{i}}|+|F_{t_{i}}(Y)-\widetilde{F}_{t_{i}}(\widetilde{Y})|+|y^{\prime}_{t_{i}}-\tilde{y}^{\prime}_{t_{i}}|$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\widetilde{F}_{0}(\tilde{y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\widetilde{X},p}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p},$
@@
$\|Y^{\prime}-\widetilde{Y}^{\prime}\|_{p,[t_{i},t_{i+1}]}+\|R^{Y}-R^{\widetilde{Y}}\|_{\frac{p}{2},[t_{i},t_{i+1}]}$
$\quad\lesssim|y_{0}-\tilde{y}_{0}|+|F_{0}(y)-\widetilde{F}_{0}(\tilde{y})|+\|y,y^{\prime};\tilde{y},\tilde{y}^{\prime}\|_{X,\widetilde{X},p}+C_{F-\widetilde{F}}+\|\mathbf{X};\widetilde{\mathbf{X}}\|_{p}.$